OpenStudy (anonymous):

On Fridays at the ballpark, a cup of soda costs $3, and each refill costs $0.80. If x represents the number of refills, which of the following equations represents the amount a person spends on soda?

OpenStudy (anonymous):

Answer Choices: y = 0.8x + 3 y = 0.8x + 3.8 y = 3x + 3.8 y = 3x + 0.8

OpenStudy (anonymous):

@asnaseer I need help i dont really understand this

OpenStudy (anonymous):

4.00+0.70X=AMOUNT SPENT BASED ON THE NUMBER OF REFILLS (X)